find three consecutive even integters such that five times the second equals twice the sum of the first and third
Found 2 solutions by mathispowerful, checkley77:
Answer by mathispowerful(115)

find three consecutive even integters such that five times the second equals twice the sum of the first and third
Let the three integers be: x, x+2, x+4
then 5(x+2) = 2[x+(x+4)]
solve it: 5x + 10 = 2(2x + 4)
simplify 5x + 10 = 4x + 8
5x -4x = 8 - 10
x = -2
so the three integers are -2, 0, 2

Answer by checkley77(12844)

x, x+2, x+4 are the 3 integers.
5(x+2)=2(x+x+4)
5x+10=2(2x+4)
5x+10=4x+8
5x-4x=8-10
x=-2 the first integer.
-2+2=0 for the second integer.
-2+4=2 for the largest integer.
Proof:
5(0)=2(-2+2)
0=2*0
0=0